OpenStudy (anonymous):

what is the answer to 5/8 + 1/4 =?/8 + ?/8 = ?/?

OpenStudy (anonymous):

\[\frac{5}{8}+\frac{1}{4}=\frac{5}{8}+\frac{2}{8}=\frac{7}{8}\]